# LTOS: Layout-controllable Text-object Synthesis via Adaptive Cross-attention Fusions

### v3

@@ II-B Dataset Construction
The dataset construction involves three steps. Step 1: Select regions suitable for text rendering by using texture and depth information of images; Step 2: Set filtering rules to generate text as required and color the text based on Color-model [gupta2016synthetic]; Step 3: Perform Poisson image editing [perez2023poisson] to blend the text into the scene. The filtering rules in step 2 include: 1) Text region number ranges from 1 to 8; 2) The number of text lines ranges from 1 to 8; 3) The text is generated in random regions that is suitable for image’s text rendering; 4) Each text region exceeds 2% of the image in length and width, with a minimum area of 5% of the entire image; 5) Text is randomly rotated, twisted, bolded, and bordered. Following that workflow, we notice that the background of text regions recognized as suitable for text rendering was too cluttered. To address the issues, we first replaced the original depth maps via recent Zoedepth [bhat2023zoedepth] to estimate more precise depth; then, we applied PPOCRv3 [li2022pp] to filter out all the images whose textual content is not satisfactory, i.e. containing low-quality renderings. These above designing enables stable and higher-quality text rendering, comparable to the real dataset in terms of depth, texture, and geometry.
LTOS dataset includes 228k samples. Each sample has: 1) an image with rendered text, 2) a caption of the image used as the prompt, 3) grounded word-level text annotations with enough information to obtain the corresponding glyph image, and 4) grounded object annotations. LTOS has some fine-grained attributes: depth of the visual text, multiple thicknesses, multiple text regions, and different rotation angles.
@@ III Experiment
Main results. As Tb I, our model significantly outperform all state-of-the-art baselines on the benchmarks in all metrics (+7.10% on OCR. ACC and +6.63% on NED). TextDiffuser generates text layout which is not effective at rendering non-center text. When AnyText renders multiple texts with different fonts, sizes and rotation angles on backgrounds with different textures, text’s clarity is affected.
Notice that GlyphControl performs poorly when rendering rotated text beyond 20 degrees especially on complex background textures.
@@
Qualitative results. Our qualitative results compared with baselines and widely used T2I generation APIs (SD-XL, Midjourney, DeepFloyd IF, and Bing Image Creator) are visualized in Fig. 3 (a). It shows that our method renders more legible, clearer visual texts and plausible objects at the predefined positions. Our model supports arbitrary rotation of text area, tiny font, and different thicknesses of the font in Fig.  3 (b). We observed limitations in baselines. AnyText can’t easily wrap and rotate text at special angles resulting in garbled text. TextDiffuser and GlyphControl fail to render tiny-font text and some special characters. They all sometimes generate objects of the wrong categories or omit objects.

### v4

@@ II-B Dataset Construction
The dataset construction involves three steps. Step 1: Select regions suitable for text rendering by using texture and depth information of images; Step 2: Set filtering rules to generate text as required and color the text based on Color-model [gupta2016synthetic]; Step 3: Perform Poisson image editing [perez2023poisson] to blend the text into the scene. The filtering rules in step 2 include: 1) Text region number ranges from 1 to 8; 2) The number of text lines ranges from 1 to 8; 3) The text is generated in random regions that is suitable for image’s text rendering; 4) Each text region exceeds 2% of the image in length and width, with a minimum area of 5% of the entire image; 5) Text is randomly rotated, twisted, bolded, and bordered.
Following that workflow, we notice that the background of text regions recognized as suitable for text rendering was too cluttered. To address the issues, we first replaced the original depth maps via recent Zoedepth [bhat2023zoedepth] to estimate more precise depth; then, we applied PPOCRv3 [li2022pp] to filter out all the images whose textual content is not satisfactory, i.e. containing low-quality renderings. These above designing enables stable and higher-quality text rendering, comparable to the real dataset in terms of depth, texture, and geometry.
LTOS dataset includes 228k samples. Each sample has: 1) an image with rendered text, 2) a caption of the image used as the prompt, 3) grounded word-level text annotations with enough information to obtain the corresponding glyph image, and 4) grounded object annotations. LTOS has some fine-grained attributes: depth of the visual text, multiple thicknesses, multiple text regions, and different rotation angles. As shown in the Fig  3, an example of our LTOS dataset.
Figure 3: An example sample of our LTOS dataset.
@@ III Experiment
Main results. As Tb I, our model significantly outperform all state-of-the-art baselines on the benchmarks in all metrics (+7.10% on OCR. ACC and +6.63% on NED). TextDiffuser generates text layout which is not effective at rendering non-center text. When AnyText renders multiple texts with different fonts, sizes and rotation angles on backgrounds with different textures, text’s clarity is affected. Notice that GlyphControl performs poorly when rendering rotated text beyond 20 degrees especially on complex background textures. Even after fine-tuning on LOST dataset, results are not ideal.
@@
Figure 4: Qualitative comparison of TOF with the baselines and widely used T2I generation APIs.
@@
Qualitative results. Our qualitative results compared with baselines and widely used T2I generation APIs (SD-XL, Midjourney, DeepFloyd IF, and Bing Image Creator) are visualized in Fig. 4. It shows that our method renders more legible, clearer visual texts and plausible objects at the predefined positions. Our model supports arbitrary rotation of text area, tiny font, and different thicknesses of the font in Fig. 5. We observed limitations in baselines. AnyText can’t easily wrap and rotate text at special angles resulting in garbled text. TextDiffuser and GlyphControl fail to render tiny-font text and some special characters. They all sometimes generate objects of the wrong categories or omit objects.
@@ IV CONCLUSION
V Appendix
LTOS TASK WITHOUT OBJECT LAYOUT. In Fig. 6, we further show several examples generated without the object layout (empty layout map as input for inference). Compared with the baselines, our model maintains a plausible generation ability even after removing the object layout, with clear multi-line text rendered in harmony with the image content.
Figure 6: Qualitative results without object layout during inference.
LTOS TASK WITH ONLY TEXT PROMPT To demonstrate the ability of TOF to generate objects, we generate images without glyph images and object layout maps and show results in Fig. 7,. The results show that even without object control, the corresponding objects can be properly generated through only text prompts, such as the blue jay and the brown wooden table. It demonstrates that TOF maintains the ability to generate plausible objects.
Figure 7: Qualitative results with only text prompt.